OpenStudy (anonymous):

what is value of x and y? x+12y=17,x-y=-9

OpenStudy (anonymous):

x=-7,y=2 @Blu-Girl

Nnesha (nnesha):

solve one equation for x or y and then plug the value into the other equation

OpenStudy (michele_laino):

Like before, please, try using the substitution method, for example: \[x=-9+y\] and substitute it in the first equation

Nnesha (nnesha):

 solve this equation for x by add y both side

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

http://openstudy.com/updates/546268fee4b05b0e1f172220 fan+medal help attatchment and hope u help

OpenStudy (anonymous):

x=-8+y,-x-6y=29 , so from that one is should be able to figure what this is

OpenStudy (anonymous):

thts ok

OpenStudy (jhannybean):

use the elimination method since you have 2 equations. x+12y = 17 x-y =-9 If we eliminate one variable, it will be easy to solve for the other. Lets multiply 12 to the second equation x+ 12y = 17 12x -12y = -108 now we add whats remaining 13x +0y = -91

OpenStudy (jhannybean):

x= -91/13 = -7 -7 - y = -9 -y = -9 + 7 y = 2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and how did u get x

OpenStudy (jhannybean):

After I eliminated y by multiplying my second equation by 12

OpenStudy (jhannybean):

Look up a few posts from this one, I explained how to do it there :)

OpenStudy (anonymous):

ok I will thx

OpenStudy (jhannybean):

13x = -91 x = -91/13

OpenStudy (anonymous):

ok thx

OpenStudy (jhannybean):

No problem :)

OpenStudy (michele_laino):

after the substitution which I have wrote, you can find y=2